OpenStudy (anonymous):

Simplify: (3y)^4 * y^-2

OpenStudy (luigi0210):

When something is raised to a power, multiply the powers. And a negative exponent, flip it. \[\Huge (a^x)^y=a^{x*y}\] \[\LARGE x^{-a}=\frac{1}{x^a}\]

OpenStudy (anonymous):

I don't really understand what you're saying.

OpenStudy (whpalmer4):

\[(3y)^4*y^{-2}\]Let's work the \((3y)^4\) part first. Do you know that \((ab)^n = a^n*b^n\)?

OpenStudy (luigi0210):

\[\LARGE (3y)^4=3^{1*4}(y^{1*4})\]

OpenStudy (anonymous):

So it would be \[3^{4}*y ^{4}\] for the first part

OpenStudy (whpalmer4):

Yes. And when you multiply two items having exponents and the same base, such as \(y^4*y^{-2}\) you keep the base and add the exponents: \(a^m*a^n = a^{m+n}\)

OpenStudy (anonymous):

Okay so what should I do next??

OpenStudy (anonymous):

Okay I'm pretty sure the answer is 3y^2. Can anyone verify this?

OpenStudy (whpalmer4):

No, it's not. \(3^4*y^4*y^{-2} = 3^4*y^{4-2} = \)

OpenStudy (luigi0210):

My mistake, sorry xD

OpenStudy (whpalmer4):

@lasoftballchic101 what's your new, corrected answer?